OpenStudy (anonymous):

area of a quadrilateral with 4 sides that's length are 5,1,4, and 4

OpenStudy (saifoo.khan):

Something is missing for sure!

OpenStudy (anonymous):

divide it into triangles and sum the areas up

OpenStudy (anonymous):



OpenStudy (anonymous):

thats what the shape looks like

OpenStudy (anonymous):

divide it into a rectangle and a triangle and add the areas

OpenStudy (anonymous):



OpenStudy (anonymous):

the area of the rectangle is 1 x4 =4 and the area of the traingle is 1/2 base x height = 1/2 x 3 x 4=6 therfore the total area is 10 units^2

OpenStudy (anonymous):

ooooooh i get it. thanks!

OpenStudy (saifoo.khan):

RIGHT^